2) Youve obtained the following balance sheet information for 2015 & 2016 along with income statement information for 2016 Catego 2015 $50 2016 Cate 2016 $50 $80 Sales Cost of Goods Sold Interest Expense Current Assets 420 60 $24 $5 Net Fixed Assets Accounts Payable Accrued Expenses 250 30 25 150 95 40 250 150 $6 30% Depreciation Long-term Debt Owners Equity Tax Rate a)Calculate the Operating Cash Flow for 2016 b)Calculate the Free Cash Flow for 2016 c) What is the free cash flow to creditors? What is the free cash flow to bondholders? d)If accounts payable increases, briefly explain in a sentence or two, how that would affect free cash flow for the company

Please refrain from using excel thanks!
Answer #1

EBIT=sales-cost of goods sold-depreciaiton
=50-24-6=20
Tax= (ebit-interest)*tax
=(20-5)*30%=4.5
a)operating cash flow=EBIT+depreciation-taxes
operating cash flow=20+6-4.5=21.50
b)Free cash flow for 2016=OCF – Change in Net working capital – Net capital spending
Change in Net working capital=(current asset end-current asset beg)-(current liab end-current liab beg)
Here current laibilites are account payble and accrued expenses
=(80-50)-((60+40)-(30+25))=-15
Net capital spending = Depreciation + Increase in fixed assets
=6+(420-250)=176
Free cash fow=21.50-(-15)-176
=-139.50
c)Cash flow to creditors = Interest – net borrowing
Net borrowing means =(long term debt end-long term debt beg)
Cash flow to creditors=5-(250-150)
=-95
free cash flow to bondholders is nothing but the interest expense and it is 5
d) The free cash flow to company will increase if account payable increase as it is like credit given to us by supplier and it increases the free cash flow to firm